OpenStudy (anonymous):

(Newton's Law of Motion) What is the average force exerted by a shot putter on a 7kg shot if the shot is moved through the distance of 2.8m and is released with a speed of 13m/s?

OpenStudy (espex):

Acceleration is assumed constant so list off your known and unknown: \[m=7kg,\space x_i=0,\space x_f=2.8m, \space v_f=\frac{13m}{s}\] Want to know propelling force \[F_p=?\] \[v=\frac{\Delta x}{t} \rightarrow \frac{v}{\Delta x} \rightarrow \frac{13m/s}{2.8m}=4.64 \sec\] \[v_f=v_i +at \rightarrow \frac{v_f-v_i}{t}=a \rightarrow \frac{13m/s - 0}{4.64\sec}=2.80\frac{m}{s^2}\] \[F_p=ma \rightarrow 4.64\frac{m}{s^2} \times 7kg=32.48\frac{kg*m}{s^2} \rightarrow 32.48N\] As always, double check my math.